What is the final temperature of a 100. g sample of iron when it loses 2.0 kJ of heat if the initial temperature was 80.0 °C?
1
Expert's answer
2020-05-24T02:37:09-0400

According to the heat equation:

Q = m × c × ∆T = m × c × (T2 - T1)

where Q - amount of heat, m - mass, c - specific heat, ∆T - change in temperature.

From here:

T2 = [Q / (m × c)] + T1

As ciron = 0.450 J/g °C:

T2 = [- 2.0 kJ / (100.0 g × 0.450 J/g °C)] + 80.0 °C) = 36.0 °C


Answer: 36.0 °C